OpenStudy (anonymous):

It takes 15 minutes to walk to the mall from your house at an average velocity of 1.1 m/s to the east. What is the displacement from your house to the mall?

jigglypuff314 (jigglypuff314):

displacement = V(average) * time and units have to match 15min = 15 * 60 = 900sec 1.1 * 900 = 990 so 990 meters East